OpenStudy (anonymous):

What is the slope of the graph of 8x – 4y = 11?

Parth (parthkohli):

First subtract both sides by 8x, then divide both sides by -4. You'll get a form like \(y = mx + b\), where m will be the slope.

OpenStudy (anonymous):

Y=mx+c m=slope ,put the equation in that form first and get the slope

OpenStudy (anonymous):

8x – 4y = 11 -4y = -8x + 11 y = 2x - 11/4 y = mx + b m = slope b = y-intercept Therefore, m = 2

OpenStudy (anonymous):

ohhh ok i get it!!!! :) everyone on this site is actually helping me understand math better lol thanks!

OpenStudy (anonymous):

You're welcome! :)